OpenStudy (anonymous):

16w+4x=y solve for x

OpenStudy (anonymous):

4x=y-16w divide both sides with 4 you will have x

OpenStudy (anonymous):

4x=-16w+y x=-4w+(1/4)y

OpenStudy (aegina):

16w+4x=y 4x=y-16w x=(1/4)y-4w x= (y-16w)/4